Final state distributions upon vibrational predissociation of the HCl dimer

\underline{L. Oudejans} and R. E. Miller, Department of Chemistry, University of North Carolina, Chapel Hill, NC 27599.

Photofragment angular distributions have been obtained for the vibrational predissociation of the HCl dimer. The initial state was selected by exciting the ”free”- or ”bounded” H-Cl stretch with an F-center laser. Assignment of the final state distributions allows for accurate determination of the dissociation energy of the dimer to be 439(2) cm^{−1}, in good agreement with the recent study of Ni etal ^{a}. The influence of the hyperfine interaction on the angular distributions will be discussed.

aH. Ni, J. M. Serafin and J. J. Valentini J. Chem. Phys. \underline{104} 2259 (1996)